OpenStudy (jiteshmeghwal9):

If \(\LARGE{a^{1/2}+b^{1/2}-c^{1/2}=0}\) then the value of \(\LARGE{(a+b-c)^2=?}\)

OpenStudy (jiteshmeghwal9):

Plz help :)

OpenStudy (anonymous):

a^1/2 + b^1/2 -c^1/2 =0 a^1/2 + b^1/2 =c^1/2 Squaring both sides gives, a+2(ab)^1/2 + b =c a+b-c = 2 (ab)^1/2 So, (a+b-c)^2 = { 2 (ab)^1/2 }^2 = 4ab

OpenStudy (jiteshmeghwal9):

\(a+b-c=-2(ab)^{1/2}\)

OpenStudy (jiteshmeghwal9):

Well...I gt it thanx a lot dude :)

OpenStudy (anonymous):

oh yes.... But the final answer is same

OpenStudy (jiteshmeghwal9):

^_^